i am absolutly confused on this question ive tried several different angles:
5x+2y+z=-11
2x-3y-z= 17
7x+y2+z= -4

Answer by ilana(307)

You have 3 equations, call them A, B, and C. First, combine A and B, eliminating z, to get an equation with x and y. Then combine B and C, again eliminating z, to get another equation with x and y. Then you will have a system with 2 equations and 2 variables, which you should know how to solve already.
Adding A and B gives 7x-y=6 (call this D). Adding B and C gives 9x-y=13 (call this E). Change equation D to y=7x-6. Now plug that y into E to get 9x-(7x-6)=13, so 2x+6=13, so 2x=7, so x=3.5. plug that back into D to get y=7(3.5)-6=24.5-6 =18.5. Using equation B, z=2x-3y-17=2(3.5)-3(18.5)-17= 7-55.5-17= -48.5-17=-65.5
So (x, y, z)=(3.5, 18.5, -65.5).